PLTOTF(1) PLTOTF(1) NAME pltotf-convertpropertylistfilestofontmetric(tfm)format SYNOPSIS pltotf[-verbose]plfilenametfmfilename DESCRIPTION Thepltotfprogramtranslatesa(human-oriented)propertylistfiletoa(program-oriented)font metricfile.Thus,aftereditingthepropertylistfile,aTFMfilecanbegeneratedforusewith,for example,tex(1). Boththeplfilenameandthetfmfilenamemustbecomplete.Noaddingofdefaultextensions orpathsearchingisdone. OPTIONS Withoutthe-verboseoption,pltotfoperatessilently.Withit,abannerandprogressreportare printedonstdout.Seetex(1)forthedetailsofspecifyingoptions. SEEALSO tftopl(1) DonaldKnuth,TeXware,TheMETAFONTbook AUTHORS DonaldKnuthwrotetheprogram.ItwaspublishedaspartoftheTeXwaretechnicalreport,availablefromtheTeXUsersGroup.HowardTrickey,PavelCurtis,andRichardFurutaoriginally portedittoUnix. 7Jan92 1